what is the equation for a line with a slope of 1/2 and passes through the point (-2,3)
Answer by checkley75(3666)

Y=mX+b
3=(1/2)*-2+b
3=-1+b
b=3+1
b=4 OR THE EQUATION IS:
Y=X/2+4